find sin[cot -1 {cos(tan -1 x)}]

find
sin[cot-1{cos(tan-1x)}]  

Grade:

1 Answers

Havishma Haranath
9 years ago
tan-1x=k then tan k =x then cos k=1/(1+x2)1/2
now sin(cot-1 1/(1+x2)1/2)
keep cot-11/(1+x2)1/2  = m and cotm = 1/(1+x2)1/2 then sinm=(1+x)1/2/(2+x2)1/2